# Draw with Thought: Unleashing Multimodal Reasoning for Scientific Diagram Generation

### v1

@@ 4.1 Coarse-to-Fine Planning for Structured Visual Understanding
To reduce perceptual complexity, we first induce a symbolic abstraction $T_{percept}$ that encodes diagram structure through layout grouping, hierarchy, visual encoding, and topology extraction. Inspired by GoT [16], we prompt Multimodal Large language Model $\mathcal{M}_{MLLM}$ to simulate perceptual reasoning via ${Thought}_{perceptual}$:
$T_{percept}\sim\mathcal{M}_{MLLM}(\cdot\mid D_{scientific},{Thought}_{percept})$ (1)
@@
$T_{percept}=(T_{gestalt},T_{hierarchy},T_{encoding},T_{connector})$
@@
$T_{hierarchy}\sim\mathcal{M}_{MLLM}(\cdot\mid T_{percept},{Thought}_{hierarchy})$ (2)
@@
$T_{hierarchy}=(\mathcal{R},\mathcal{E},\mathcal{L})$ (3)
@@
=$\mathcal{L}=\{\texttt{align}(e_{i},e_{j}),\ \texttt{connect}(e_{i}\rightarrow e_{j}),\ \texttt{layer}(e_{k},z)\}$ (4)
@@
4.2 Structure-Aware Code Generation via Progressive Realization
@@
Conditioned on $T_{hierarchy}$, we prompt $\mathcal{M}_{MLLM}$ to progressively synthesize the final XML code $Y_{mxGraph}$:
$Y_{mxGraph}\sim\mathcal{M}_{MLLM}(\cdot\mid T_{hierarchy},{Thought}_{code})$ (5)
@@
$Y^{(0)}_{mxGraph}=(Y_{doc},Y_{style},Y_{node},Y_{layout},Y_{edge})$ (6)
@@
$Y^{(0)}_{mxGraph}$ $=\bigcup_{i}Y_{i},$ (7)
for $\quad\text{for }Y_{i}\in\{Y_{doc},Y_{style},Y_{node},Y_{layout},Y_{edge}\}$
Each component serves a distinct role: $Y_{doc}$ initializes the document tree (e.g., mxfile, diagram); $Y_{style}$ defines reusable style tokens (e.g., color, font, stroke); $Y_{node}$ instantiates entities from $\mathcal{E}$ with geometry and ID; $Y_{layout}$ applies spatial constraints from $\mathcal{L}$; and $Y_{edge}$ encodes directed connectors with routing semantics.
@@
While structurally grounded, initial outputs often contain XML formatting flaws—such as unclosed tags, incorrect nesting, or inconsistent styles for long-context for output during inference [7]. To ensure syntactic correctness and structural executability, we introduce a multi-round refinement module guided by format-level reasoning.
At each refinement step $t$, $\mathcal{M}_{MLLM}$ revisits the prior output $Y^{(t-1)}_{mxGraph}$ and produces an updated version using format-specific reasoning via ${RefineThought}^{(t)}_{format}$ prompting:
$Y^{(t)}_{mxGraph}\sim$ $\mathcal{M}_{MLLM}\big{(}Y^{(t-1)}_{mxGraph},\ {RefineThought}^{(t)}_{format}\big{)},$
$\quad t=1,\dots,T_{refine}$ (8)
@@
if $\text{if }\texttt{Draw.ioVerifier}(Y^{(t)}_{\text{mxGraph}})=\texttt{valid}\text{ then break}$ (9)
@@
[…]
@@
5 Experiment
5.1 Experimental Setup
To provide a comprehensive evaluation, we compared Draw-with-Thought (DwT) against eight state-of-the-art MLLMs— GPT-4o  [^2] We use the ”gpt4o-2024-07-18” version of API for reproducibility. [39], Claude3.5-sonnet [4], Claude3.7-sonnet [5], Gemini-2.0 [23], Grok-3 [59], Llama-3.2V-11B[2] and Qwen2.5-VL-32B [6].
Following [44, 61], we evaluate performance across two key dimensions using four metrics: (1) Similarity, measured by CLIP Score [26, 47], DINO Score [11, 40], and FID [27], to assess semantic and visual alignment with ground truth diagrams; (2) Aesthetics, assessed using the Aesthetic Score [45], which reflects perceived visual quality. Representative results are shown in Figure 3, with full visualizations and quantitative comparisons provided in Appendix C.2.
@@ 5.1 Experimental Setup
Our method achieves remarkable improvements, particularly in Hard difficulty tasks, by fundamentally aligning with the cognitive processes underlying diagram comprehension[1, 3]. By guiding models through visual decomposition, structural inference, and symbolic code synthesis, our approach directly addresses the core challenges in scientific diagram parsing: maintaining coherent spatial-relational representations while translating between visual and symbolic modalities [58]. The exceptional improvement in Hard tasks demonstrates that our method effectively scaffolds the complex reasoning required to handle diagrams with intricate spatial arrangements, nested hierarchies, and complex logical relationships—precisely the elements that make scientific diagrams challenging to reconstruct.
@@
The consistent performance degradation observed across all models as task difficulty increases directly reflects the escalating cognitive demands of complex diagram parsing. This degradation is particularly pronounced in base models without reasoning scaffolds, aligning with investigations of MLLM cognitive and memory mechanisms [41], which indicate that complex visual-spatial tasks can overwhelm working memory when approached holistically [49, 36]. The disproportionate decline in FID scores compared to semantic alignment metrics suggests that models primarily struggle with maintaining spatial coherence and structural relationships rather than conceptual understanding.
@@
Current MLLMs face a fundamental bottleneck in scientific diagram reconstruction: capturing multi-level abstractions often requires generating outputs exceeding the model’s token limit, leading to truncation or content loss. While models like Claude 3.7-sonnet benefits from strong cross-modal consistency, its sharp performance drop on complex diagrams—reflected by a 26.8% decline in CLIP score and a 105.5% increase in FID from Easy to Hard tasks (Table 2)—reveals the inadequacy of implicit spatial reasoning [63]. Conversely, the 47.2% CLIP score drop observed in GPT-4o suggests that certain architectures fundamentally lack the capacity to preserve coherent spatial-relational representations across varying levels of diagrammatic complexity. Given that spatial arrangements encode underlying logical relationships, the inability to maintain such structure compromises semantic fidelity and highlights spatial-relational understanding as a critical yet underdeveloped competency in existing multimodal models.
Figure 5: Comparison of diagram reconstruction results across different approaches. (a) Raw image shows the original scientific diagram with a two-step process structure. (b) Claude-3.7-sonnet’s direct generation attempt shows reasonable structure but misses some key connections and node relationships. (c) Our Draw with Thought (DwT) approach produces a more accurate reconstruction that better preserves the hierarchical structure, node positions, and connection relationships from the original diagram.
5.2 Ablation Study
Our Draw with Thought consists of three key components:—Perceptual Structuring, Semantic Layout Planning, and Structured Code Generation—which collectively support the extraction, organization, and formalization of visual content from scientific diagrams. To further enhance structural reliability, our framework incorporates a hierarchical XML generation module coupled with a self-refinement mechanism that iteratively corrects syntactic inconsistencies.
As shown in Table 3, we conduct ablation studies on our Plot2XML using LLaMA 3.2-V-11B, evaluating CLIP and DINO for semantic alignment, aesthetic score for visual quality, and XML Validation for structural correctness (Table 3). The full model yields the best results (CLIP 0.657, DINO 0.815, Aesthetics 3.807, Validation 89%). Removing Perceptual Structuring reduces CLIP and DINO by 7.2% and 9.0%, respectively, indicating weaker semantic grounding, while excluding Layout Planning leads to the largest overall decline—CLIP drops 25.1%, aesthetics 7.0%, and validation falls to 73%—highlighting its role in spatial-structural coherence. Excluding Context Style Design causes moderate losses in CLIP (–11.4%) and aesthetics (–7.0%), suggesting its effect is primarily stylistic. Removing Hierarchical XML generation lowers validation to 66%, showing flat generation alone is insufficient, while refinement partially recovers it to 84%, underscoring the value of structural priors. These results demonstrate that each module contributes distinct functional value, and that high-fidelity, executable diagrams rely on the synergy of perceptual grounding, spatial reasoning, and structure-aware generation.
Table 3: Ablation experiment. Results for Draw with Thought and ablation with different components using LLaMA 3.2-V-11B on Plot2XML.
@@ 5.2 Ablation Study
5.3 Human Evaluation
Following [30, 35, 18], we conducted a human evaluation to assess the alignment between automated metrics and human perception, combining absolute ratings and comparative judgments. For absolute ratings, domain experts used a 10-point scale [30] to evaluate each reconstructed diagram in terms of similarity to the original (1 = “completely different” to 10 = “identical”) and aesthetic quality (1 = “poor quality” to 10 = “professional quality”). For comparative evaluation, we employed Best-Worst Scaling (BWS) [35, 18], where annotators selected the best and worst outputs within a presented set. To manage annotation effort, comparisons were limited to the two strongest baseline models—Qwen2.5-VL and Claude 3.7-Sonnet—alongside our method (DWT), with human-annotated diagrams serving as reference.
As shown in Table 4, DWT achieves the highest scores in both similarity (7.3) and aesthetics (8.1), closely approaching the human reference (9.4 / 9.0), and outperforming Claude (6.1 / 7.0) and Qwen (5.6 / 6.3). In BWS, DWT receives the highest score (0.74), indicating consistent preference, while Qwen’s negative score (–0.87) reflects frequent disfavor. These results confirm DWT’s superior semantic fidelity, visual quality, and alignment with human preferences. Annotation protocol details for human evaluation are provided in Appendix B.
Table 4: Human Evaluation Results. DWT surpasses all automated methods and closely matches human-created diagrams across all dimensions.
@@ 5.3 Human Evaluation
Spearman correlation between model metrics and human evaluations [38] shows that CLIP scores exhibit the strongest correlation with human judgments of similarity (Spearman’s $\rho=0.825$), followed by DINO ($\rho=0.763$) and FID ($\rho=-0.712$). This strong correlation with CLIP scores is particularly noteworthy, as it suggests that semantic alignment between the original and reconstructed diagrams—precisely what CLIP scores measures—is the most salient factor in human perception of reconstruction quality. The substantial negative correlation with FID scores indicates that while low-level visual fidelity matters, humans prioritize semantic and structural coherence over pixel-perfect reproduction when evaluating scientific diagrams.
6 Conclusion
@@
In this paper, we present Draw with Thought, a training-free framework unleashing Multimodal Large Language Models (MLLMs) for reconstructing scientific diagrams from rasterized images into editable, executable mxGraph XML. Grounded in cognitive load [41] and structure mapping [22] theory, our Draw with Thought involving Coarse-to-Fine Planning for Structured Visual Understanding and Structure-Aware Code Generation via Progressive Realization—yields diagram representations that are both semantically faithful and structurally valid. To support systematic evaluation, we introduce Plot2XML, a benchmark of 247 real-world scientific diagrams with gold-standard XML annotations and multi-dimensional complexity analysis. We benchmarked our method across multiple state-of-the-art MLLMs, consistently demonstrating superior performance on high-complexity scientific diagrams through enhanced semantic alignment, layout fidelity, and XML validity.

### v2

@@ 4.1. Coarse-to-Fine Planning for Structured Visual Understanding
To reduce perceptual complexity, we first induce a symbolic abstraction $T_{percept}$ that encodes diagram structure through layout grouping, hierarchy, visual encoding, and topology extraction. Inspired by GoT (Fang et al., 2025), we prompt Multimodal Large language Model $\mathcal{M}_{MLLM}$ to simulate perceptual reasoning via ${Thought}_{perceptual}$:
(1) $T_{percept}\sim\mathcal{M}_{MLLM}(\cdot\mid D_{scientific},{Thought}_{percept}),$
@@
(2) $T_{percept}=(T_{gestalt},T_{hierarchy},T_{encoding},T_{connector}).$
@@
(3) $T_{hierarchy}\sim\mathcal{M}_{MLLM}(\cdot\mid T_{percept},{Thought}_{hierarchy}),$
@@
(4) $T_{hierarchy}=(\mathcal{R},\mathcal{E},\mathcal{L}),$
@@
(5) $\mathcal{L}=\{\texttt{align}(e_{i},e_{j}),\ \texttt{connect}(e_{i}\rightarrow e_{j}),\ \texttt{layer}(e_{k},z)\}.$
@@
4.2. Structure-Aware Code Generation via Progressive Realization
@@
Based on $T_{hierarchy}$, we prompt $\mathcal{M}_{MLLM}$ to progressively synthesize the final XML code $Y_{mxGraph}$:
(6) $Y_{mxGraph}\sim\mathcal{M}_{MLLM}(\cdot\mid T_{hierarchy},{Thought}_{code}),$
@@
(7) $Y^{(0)}_{mxGraph}=(Y_{doc},Y_{style},Y_{node},Y_{layout},Y_{edge}),$
@@
(8) $Y^{(0)}_{mxGraph}$ $=\bigcup_{i}Y_{i},$
$\quad\text{for }Y_{i}\in\{Y_{doc},Y_{style},Y_{node},Y_{layout},Y_{edge}\}.$
Figure 4. Qualitative Comparison Between DwT and State-of-the-Art MLLMs. We visualize the output results. We observe that our model can understand the layout and the content of the image well, achieving better performance.
Each component serves a distinct role: $Y_{doc}$ initializes the document tree (e.g., mxfile, diagram); $Y_{style}$ defines reusable style tokens (e.g., color, font, stroke); $Y_{node}$ instantiates entities $\mathcal{E}$ with geometry and ID; $Y_{layout}$ applies spatial constraints from $\mathcal{L}$; and $Y_{edge}$ encodes directed connectors with routing semantics.
@@
While structurally grounded, initial outputs often contain XML formatting flaws—such as unclosed tags, incorrect nesting, or inconsistent styles for long-context for output during inference (Bai et al., 2024). To ensure syntactic correctness and structural executability, we introduce a multi-round refinement module guided by format-level reasoning.
At each refinement step $t$, $\mathcal{M}_{MLLM}$ revisit the prior output $Y^{(t-1)}_{mxGraph}$ and produce an updated version using format-specific reasoning via ${RefineThought}^{(t)}_{format}$ prompting:
$Y^{(t)}_{mxGraph}\sim\$$\mathcal{M}_{MLLM}\big(Y^{(t-1)}_{mxGraph},\ {RefineThought}^{(t)}_{format}\big)$,
(9) $\quad t=1,\dots,T_{refine}$
@@
(10) $\text{if }\texttt{Draw.ioVerifier}(Y^{(t)}_{\text{mxGraph}})=\texttt{valid}\text{ then break}.$
@@
Table 2. Comprehensive performance comparison of AI models across three difficulty levels. Our visual design employs a gradient blue background (increasing intensity with difficulty) to enhance readability, with purple section headers and a teal highlight for the best-performing model. Star icons (˘, ˇ) indicate difficulty levels, while specialized icons represent model categories: ?for base models, for Chain-of-Thought models, and for our proposed method. Performance metrics include CLIP and DINO scores measuring semantic consistency, FID score for image quality, aesthetic rating, and token consumption in thousands (K).
Model Easy Difficulty ˘ Medium Difficulty ˇ˘ Hard Difficulty ˇˇˇ
@@
5. Experiment
5.1. Experimental Setup
To provide a comprehensive evaluation, we compared Draw-with-Thought (DwT) against eight state-of-the-art MLLMs— GPT-4o (OpenAI, 2024), Claude3.5-sonnet (Anthropic, 2024), Claude3.7-sonnet (Anthropic, 2025), Gemini-2.0 (Google, 2024), Grok-3 (xAI, 2025), Llama-3.2V-11B(AI, 2024) and Qwen2.5-VL-32B (Bai et al., 2025).
Following (Rodriguez et al., 2023; Xing et al., 2024), we evaluate performance across two key dimensions using four metrics: (1) Similarity, measured by CLIP Score (Hessel et al., 2021; Song et al., 2022), DINO Score (Caron et al., 2021; Oquab et al., 2023), and FID (Heusel et al., 2017), to assess semantic and visual alignment with ground truth diagrams; (2) Aesthetics, assessed using the Aesthetic Score (Schuhmann, 2022), which reflects perceived visual quality. Representative results are shown in Figure 4.
@@ 5.1. Experimental Setup
Our method achieves remarkable improvements, particularly in Hard difficulty tasks, by fundamentally aligning with the cognitive processes underlying diagram comprehension(Abrahamsen and Bechtel, 2015; Allen, 2011). By guiding models through visual decomposition, structural inference, and symbolic code synthesis, our approach directly addresses the core challenges in scientific diagram parsing: maintaining coherent spatial-relational representations while translating between visual and symbolic modalities (Wu et al., 2023). The exceptional improvement in Hard tasks demonstrates that our method effectively scaffolds the complex reasoning required to handle diagrams with intricate spatial arrangements, nested hierarchies, and complex logical relationships—precisely the elements that make scientific diagrams challenging to reconstruct.
@@
The consistent performance degradation observed across all models as task difficulty increases directly reflects the escalating cognitive demands of complex diagram parsing. This degradation is particularly pronounced in base models without reasoning scaffolds, aligning with investigations of MLLM cognitive and memory mechanisms (Plass et al., 2010), which indicate that complex visual-spatial tasks can overwhelm working memory when approached holistically (Srivastava et al., 2022; Lu et al., 2021). The disproportionate decline in FID scores compared to semantic alignment metrics suggests that models primarily struggle with maintaining spatial coherence and structural relationships rather than conceptual understanding.
@@
Current MLLMs face a fundamental bottleneck in scientific diagram reconstruction: capturing multi-level abstractions often requires generating outputs exceeding the model’s token limit, leading to truncation or content loss. While models like Claude 3.7-sonnet benefit from strong cross-modal consistency, their sharp performance drop on complex diagrams—reflected by a 26.8% decline in CLIP score and a 105.5% increase in FID from Easy to Hard tasks (Table 2)—reveals the inadequacy of implicit spatial reasoning (Ye et al., 2025). Conversely, the 47.2% CLIP score drop observed in GPT-4o suggests that certain architectures fundamentally lack the capacity to preserve coherent spatial-relational representations across varying levels of diagrammatic complexity. Given that spatial arrangements encode underlying logical relationships, the inability to maintain such structure compromises semantic fidelity and highlights spatial-relational understanding as a critical yet underdeveloped competency in existing multimodal models.
Figure 5. Comparison of diagram reconstruction results across different approaches.
Table 3. Ablation experiment. Results for Draw with Thought and ablation with different components using LLaMA 3.2-V-11B on Plot2XML.
@@
5.2. Ablation Study
Our Draw with Thought consists of three key components: Perceptual Structuring, Semantic Layout Planning, and Structured Code Generation—which collectively support the extraction, organization, and formalization of visual content from scientific diagrams. The Context Style Design refers to the style extraction and definition step from our planning phase. To further enhance structural reliability, our framework incorporates a hierarchical XML generation module coupled with a self-refinement mechanism that iteratively corrects syntactic inconsistencies. Operates within an iterative feedback loop, acting as a targeted debugging way.
As shown in Table 3, we conducted ablation studies on our Plot2XML benchmark using LLaMA 3.2-V-11B, evaluating CLIP and DINO for semantic alignment, aesthetic score for visual quality, and XML Validation for structural correctness. The results reveal a clear dependency chain, where errors in perceptual or planning stages cascade, impacting both semantic fidelity and structural integrity. This validates our design choice to separate high-level reasoning from low-level code generation, allowing for more robust and modular processing. These results demonstrate that each module contributes distinct functional value and that high-fidelity, machine-readable diagrams rely on the synergy of perceptual grounding, spatial reasoning, and structure-aware generation.
5.3. Human Evaluation
Following (Krosnick, 2017; Louviere et al., 2015; Flynn and Marley, 2014), we conducted a human evaluation to assess the alignment between automated metrics and human perception, combining absolute ratings and comparative judgments. For absolute ratings, domain experts used a 10-point scale (Krosnick, 2017) to evaluate each reconstructed diagram in terms of similarity to the original and aesthetic quality. For comparative evaluation, we employed Best-Worst Scaling (BWS) (Louviere et al., 2015; Flynn and Marley, 2014), where annotators selected the best and worst outputs within a presented set. To manage annotation effort, comparisons were limited to the two strongest baseline models—Qwen2.5-VL and Claude 3.7-Sonnet—alongside our method (DWT), with human-annotated diagrams serving as reference.
As shown in Table 4, DWT achieves the highest scores in both similarity and aesthetics, closely approaching the human reference and outperforming the best LLMs. In BWS, DWT receives the highest score, indicating consistent preference, while Qwen’s negative score reflects frequent disfavor. These results confirm DWT’s superior semantic fidelity, visual quality, and alignment with human preferences.
Table 4. Human Evaluation Results. DWT surpasses all automated methods and closely matches human-created diagrams across all dimensions.
@@ 5.3. Human Evaluation
Spearman correlation between model metrics and human evaluations (Myers and Sirois, 2014) shows that CLIP scores exhibit the strongest correlation with human judgments of similarity (Spearman’s $\rho=0.825$), followed by DINO ($\rho=0.763$) and FID ($\rho=-0.712$). This strong correlation with CLIP scores is particularly noteworthy, as it suggests that semantic alignment between the original and reconstructed diagrams—precisely what CLIP scores measure—is the most salient factor in human perception of reconstruction quality. The substantial negative correlation with FID scores indicates that while low-level visual fidelity matters, humans prioritize semantic and structural coherence over pixel-perfect reproduction when evaluating scientific diagrams.
6. Conclusion
@@
In this paper, we present Draw with Thought, a training-free framework unleashing Multimodal Large Language Models (MLLMs) for reconstructing scientific diagrams from rasterized images into editable, executable mxGraph XML. Grounded in cognitive load (Plass et al., 2010) and structure mapping (Gentner, 1983) theory, our Draw with Thought involving Coarse-to-Fine Planning for Structured Visual Understanding and Structure-Aware Code Generation via Progressive Realization—yields diagram representations that are both semantically faithful and structurally valid. To support systematic evaluation, we introduce Plot2XML, a benchmark of 247 real-world scientific diagrams with gold-standard XML annotations and multi-dimensional complexity analysis. We benchmarked our method across multiple state-of-the-art MLLMs, consistently demonstrating superior performance on high-complexity scientific diagrams through enhanced semantic alignment, layout fidelity, and XML validity.